# Context-aware Communication for Multi-agent Reinforcement Learning

### v1

@@ Abstract.
Effective communication protocols in multi-agent reinforcement learning (MARL) are critical to fostering cooperation and enhancing team performance. To leverage communication, many previous works have proposed to compress local information into a single message and broadcast it to all reachable agents. This simplistic messaging mechanism, however, may fail to provide adequate, critical, and relevant information to individual agents, especially in severely bandwidth-limited scenarios. This motivates us to develop context-aware communication schemes for MARL, aiming to deliver personalized messages to different agents. Our communication protocol, named CACOM, consists of two stages. In the first stage, agents exchange coarse representations in a broadcast fashion, providing context for the second stage. Following this, agents utilize attention mechanisms in the second stage to selectively generate messages personalized for the receivers. Furthermore, we employ the learned step size quantization (LSQ) technique for message quantization to reduce the communication overhead. To evaluate the effectiveness of CACOM, we integrate it with both actor-critic and value-based MARL algorithms. Empirical results on cooperative benchmark tasks demonstrate that CACOM provides evident performance gains over baselines under communication-constrained scenarios.
@@
Cooperative multi-agent reinforcement learning (MARL) has recently emerged as an exciting research avenue due to its applicability to real-world scenarios. Many of these applications naturally call for multi-agent frameworks, such as resource management (Ying and Dayong, 2005), package delivery (Seuken and Zilberstein, 2007), disaster rescue (Parker et al., 2016), and robots control (Swamy et al., 2020). Despite the considerable success of seminal works (Rashid et al., 2020; Lowe et al., 2017; Yu et al., 2021) in generalizing reinforcement learning algorithms to multi-agent cases under the centralized training and decentralized execution (CTDE) paradigm (Kraemer and Banerjee, 2016; Lyu et al., 2021), further improvement in MARL is hindered by a few prominent obstacles, such as non-stationarity and partial observability (Oroojlooy and Hajinezhad, 2022).
@@ 5.2. SMAC
Results: We show the performance of CACOM and baselines on SMAC benchmarks in Figure 7 and the corresponding gate pruning results in Table 2. Overall, CACOM consistently outperfoms the baselines. Compared to the MPE, the observations in SMAC are more sophisticated therefore exacerbate the consequences of limited bandwidth. As a result, most baseline communication protocols perform even worse than the communication-free method QMIX under severe bandwidth constraints (e.g., Figure 6(b) and 6(c)). In contrast, CACOM still manages to obtain evident performance gains. This highlights its ability to leverage the limited bandwidth efficiently by transmitting only the most relevant information.
@@
Ablation Studies: To verify the effectiveness of different design choices in CACOM, we conduct ablation studies on MMM3 and 27m_vs_30m scenarios, which are illustrated in Figure 6(e) and 6(f). We keep the CACOM network architecture but replace the context-aware communication with broadcasting communication and denote it as CACOM-BC. We replace the attention blocks in the network with multi-layer perceptron (MLP) blocks and denote it as CACOM-MLP. We remove the gating mechanism in CACOM and call it CACOM w/o gate. We remove the auxiliary task and call it CACOM w/o aux. Figure 6(e) and 6(f) show that CACOM outperforms CACOM-BC, indicating the better expressiveness achieved by context-aware communication. The unsatisfactory performance of CACOM-MLP implies that the success of CACOM shall also attribute to the network design. Without the attention mechanism to help identifying the related information, it is difficult for the helpers to learn to generate the personalized messages. Furthermore, the gating mechanism improves the performance by filtering out the unrelated information and the auxiliary task assists the training by stabilizing the training in some scenarios.

### v3

@@ Abstract.
Effective communication protocols in multi-agent reinforcement learning (MARL) are critical to fostering cooperation and enhancing team performance. To leverage communication, many previous works have proposed to compress local information into a single message and broadcast it to all reachable agents. This simplistic messaging mechanism, however, may fail to provide adequate, critical, and relevant information to individual agents, especially in severely bandwidth-limited scenarios. This motivates us to develop context-aware communication schemes for MARL, aiming to deliver personalized messages to different agents. Our communication protocol, named CACOM, consists of two stages. In the first stage, agents exchange coarse representations in a broadcast fashion, providing context for the second stage. Following this, agents utilize attention mechanisms in the second stage to selectively generate messages personalized for the receivers. Furthermore, we employ the learned step size quantization (LSQ) technique for message quantization to reduce the communication overhead. To evaluate the effectiveness of CACOM, we integrate it with both actor-critic and value-based MARL algorithms. Empirical results on cooperative benchmark tasks demonstrate that CACOM provides evident performance gains over baselines under communication-constrained scenarios. The code is publicly available at https://github.com/LXXXXR/CACOM.
@@
Cooperative multi-agent reinforcement learning (MARL) has recently emerged as an exciting research avenue due to its applicability to real-world scenarios. Many of these applications naturally call for multi-agent frameworks, such as resource management (Ying and Dayong, 2005), package delivery (Seuken and Zilberstein, 2007), disaster rescue (Parker et al., 2016), and robots control (Swamy et al., 2020). Despite the considerable success of seminal works (Rashid et al., 2020; Lowe et al., 2017; Yu et al., 2021) in generalizing reinforcement learning algorithms to multi-agent cases under the centralized training and decentralized execution (CTDE) paradigm (Kraemer and Banerjee, 2016; Lyu et al., 2021), further improvement in MARL is hindered by a few prominent obstacles, such as non-stationarity and partial observability (Yuan et al., 2023; Oroojlooy and Hajinezhad, 2022).
@@ 5.2. SMAC
Results: We show the performance of CACOM and baselines on SMAC benchmarks in Figure 7 and the corresponding gate pruning results in Table 2. Overall, CACOM consistently outperfoms the baselines. Compared to the MPE, the observations in SMAC are more sophisticated therefore exacerbate the consequences of limited bandwidth. As a result, most baseline communication protocols perform even worse than the communication-free method QMIX under severe bandwidth constraints (e.g., Figure 7(b) and 7(c)). In contrast, CACOM still manages to obtain evident performance gains. This highlights its ability to leverage the limited bandwidth efficiently by transmitting only the most relevant information.
@@
Ablation Studies: To verify the effectiveness of different design choices in CACOM, we conduct ablation studies on MMM3 and 27m_vs_30m scenarios, which are illustrated in Figure 7(e) and 7(f). We keep the CACOM network architecture but replace the context-aware communication with broadcasting communication and denote it as CACOM-BC. We replace the attention blocks in the network with multi-layer perceptron (MLP) blocks and denote it as CACOM-MLP. We remove the gating mechanism in CACOM and call it CACOM w/o gate. We remove the auxiliary task and call it CACOM w/o aux. Figure 7(e) and 7(f) show that CACOM outperforms CACOM-BC, indicating the better expressiveness achieved by context-aware communication. The unsatisfactory performance of CACOM-MLP implies that the success of CACOM shall also attribute to the network design. Without the attention mechanism to help identifying the related information, it is difficult for the helpers to learn to generate the personalized messages. Furthermore, the gating mechanism improves the performance by filtering out the unrelated information and the auxiliary task assists the training by stabilizing the training in some scenarios.